Calcium ions bind to the SERCA Ca^2+ -ATPase, which has two identical calcium sites, in two stages with apparent equilibrium constants K1=7×105^M^−1 and K2=2×10^6M^−1

A. Calculate K

B. Calculate τ.

C. Is the binding positively or negatively cooperative? (pick one below)

- The binding is negatively cooperative

- The binding is positively cooperative.

- The binding could be both positively and negatively cooperative.

- The binding is neither positively nor negatively cooperative.
Answer #1

1) K = K1K2 = (7*10^5)(2*10^6) = 14 *10^11 M-2

2) log(\tau) = -logK = -log (14*10^11)

\tau = 7.14*10^-13

3) since K2>K1, the binding is positively cooperative